A particle with a charge of q = -5.68nC is moving in a uniform magnetic field of B? =( 2.20T ) z^. The magnetic force on the particle is measured to beF? =( 6.10?N )?y^.

a) Calculate the x component of the velocity of the particle.

b)What is the radius of the circular motion the particle will have in the magnetic field if the particle has a mass of 0.700g ?

c)What is the period of this circular motion?

d)What is the pitch of the motion if the velocity has a y component of 4.90m/s .

Part a is -488m/s, but I can not get the answers to the other parts. Please Help!!
Answer #1

a) F=qvB
6.10\times 10^{-6}=(5.68\times 10^{-9})v(2.20)
v=488.16 m/s
b)
R=\frac{mV}{qB}
R=\frac{0.7\times 10^{-3}\times 6.10\times 10^{-6}}{5.68\times 10^{-9}(2.20)}
R=0.34 m
c)

T=\frac{2\pi m}{qB}
T=\frac{0.7\times 10^{-3}\times 2\pi }{5.68\times 10^{-9}(2.20)}